# Contextualize-then-Aggregate: Circuits for In-Context Learning in Gemma-2 2B

### v3

@@ 2 Background: Aggregation and Function Vectors
Prior work on the mechanistics of ICL in LLMs has provided a few robust insights, which are our starting point. Multiple studies have documented the existence of task vectors or function vectors (Hendel et al., 2023; Todd et al., 2024; Song et al., 2025; Kharlapenko et al., 2025; Yin and Steinhardt, 2025): Several attention heads at the last prompt token (in our case, $t_{N+1}$) output vectors that, across ICL tasks, encode task information. Erasing them prevents ICL; patching them elsewhere leads to execution of the task. Yin and Steinhardt (2025) find that Function Vector Heads are more important to ICL than Induction Heads, which had been hypothesized to be key in prior work (Olsson et al., 2022; Crosbie and Shutova, 2024) (cf. also Bansal et al. (2023)).
Todd et al. (2024) (followed by Yin and Steinhardt (2025)) operationalize Function Vector Heads as heads whose activations, when patched from valid ICL prompts to shuffled uninformative prompts, can most increase the likelihood of the desired target (i.e., better performing the task articulated by the valid prompts). Aggregating activations from these attention heads results in function vectors, which can trigger execution of the task in uninformative or zero-shot contexts. Relatedly, Kharlapenko et al. (2025) find sparse autoencoder (SAE) features that encode task information (in few-shot examples) and features that execute them (at $t_{N+1}$). However, it remains largely open what circuit assembles this information, and how information is assembled from the different few-shot examples to these function vectors. One simple hypothesis could be that each few-shot example computes a representation, and this is aggregated by Function Vector Heads.
@@
On the other hand, function vectors might not play an important role in certain kind of ICL tasks. Cho et al. (2025) suggest a somewhat different circuit: specifically, they suggest stages of induction circuits, (1) assembling an encoding of each few-shot example’s input and moving it to few-shot label position, (2) attending to labels whose input is similar to the query, and copying these labels to the query. However, their experiments are restricted to semantic text classification (e.g., sentiment), in which few-shot inputs typically are either semantically aligned or opposed. While this is a reasonable circuit, the model would presumably use other circuit with function vectors for ICL tasks involving more diversified inputs. In addition, the arguments are largely based on evaluating representational similarity and on supervised probing, which does not establish that this mechanism is causally sufficient. While a causal intervention, through zeroing out the relevant components, shows this mechanism is relevant to ICL (Section 5.1 in Cho et al. (2025)), it leaves open if the mechanism alone can fully explain the behavior, or whether it is only a part of the overall mechanism.
Kharlapenko et al. (2025) perform a causal analysis leading to a sparse feature circuit in Gemma-2 2B. However, their circuit does not distinguish between inputs and outputs for different few-shot examples, making it hard to assess whether task information is created independently at each example or whether the representations are dependent across few-shot examples.
Overall, while function vectors play an important role for many ICL tasks, it remains unclear what information flow leads to aggregation of task information across few-shot examples. To foreshadow our results, we show that, while parallel aggregation of per few-shot representations is important, the overall circuit is task-dependent, and may require a variety of additional components which, as we will show, contextualize the per few-shot representations on other few-shot examples.
@@ 3 Results
We focus on Gemma-2, with a focus on its 2B version for tractability. Our methodology allows us to also investigate larger variants (9B and 27B). We evaluated the model on tasks from Todd et al. (2024), and determined five tasks on which the 2B model achieves high (between $88\%$ and $100\%$) accuracy: Copying  [^2] This is not included in Todd et al. (2024), but we added it due to its naturalness., Present-Past, Capitalization, Country-Capital, Person-Sport (Figure 2). [^3] We also evaluated six other models at the 2B/3B scale, finding that they overall underperformed Gemma-2 2B at 10 shots on these tasks (Table 7), further motivating our focus on Gemma-2. We decided to focus on tasks with high accuracy in order to obtain a clear signal, as for these tasks there is a clear understanding of the input-output behavior exhibited by the full model. We focused experiments on 3-shot and 10-shot prompts, covering both shorter and longer prompts.
Each of the five tasks has an input space $\mathcal{X}$ (e.g., arbitrary lowercase words for Capitalization; countries for Country-Capital) and an output space $\mathcal{Y}$ (e.g., arbitrary capitalized words for Capitalization; cities for Country-Capital), and defines a function $f:\mathcal{X}\rightarrow\mathcal{Y}$ (e.g., mapping countries to their capitals) (Appendix, Figure 25). Each $N$-shot prompt then has the following structure: $x_{1}t_{1}y_{1}n_{1}\dots x_{N}t_{N}y_{N}n_{N}x_{N+1}t_{N+1}$, $\forall i:x_{i}\in\mathcal{X},y_{i}\in\mathcal{Y},y_{i}=f(x_{i}),t_{i}=\text{\textbackslash t},n_{i}=\text{\textbackslash n}$.
@@
3.1 Identifying Circuits
To localize the behavior of the model, we use patching, an approach widely adopted in the literature (Wang et al., 2022; Hanna et al., 2023). With this technique, a model is viewed as a computation graph with activations in different layers as nodes, and computations between them as edges. Then, it is possible to ablate some of the edges in the graph, forcefully replacing the computation along a specific edge with the one computed on the counterfactual input. Counterfactual inputs are designed to erase relevant information from the prompt while leaving other information intact for even better localization of model behavior. For example, when substituting "Berlin" with counterfactual "Paris", we isolate computations specific to the city’s identity, not those activated by its category (city) or token-type (word). If ablating a set of edges does not lead to a drop in the model’s performance, then the information unique to the original input relative to the counterfactual input, which was transferred along these edges, is not causal for the model’s prediction. This allows to discover the circuits - subgraphs of the full model’s computation graph that can explain a major part of model’s performance on a specific task.
In contrast to the standard approach, we focus on the information flow between positions in a prompt; consequently, we differentiate between activations of the same heads at different positions. We keep nodes for $x_{i}$ and $y_{i}$ in different examples distinct, which allows us to analyze the information flow between few-shot examples. [^4] This contrasts with the approach of Kharlapenko et al. (2025), who collapse all $x_{i}$ nodes into a single node (and same for all $y_{i}$ nodes).
We first discover a position-level circuit. Nodes are the positions in the prompt; edges are directed arcs between them, representing the flow of information (Figure 1). Importantly, in position-level circuits we collapse all edges across layers into a single edge. Ablating such an edge means ablating all corresponding edges in every attention head and layer. For illustration, $x_{2}$ in a position-level circuit can only receive four incoming edges: from $x_{1},t_{1},y_{1}$, or $n_{1}$.  [^5] The edges from/to bos/eos tokens are always ablated (Appendix E) Position-level circuits provide useful and interpretable upper bounds on information flow, though they do not distinguish between computations happening in different layers. We treated all tokens within a few-shot input $x_{i}$ or output $y_{i}$ together, as the number of tokens varies between few-shot examples. As positions, we distinguish the few-shot inputs $x_{i}$ and outputs $y_{i}$ and the separators within ($t_{i}$) and between ($n_{i}$) examples; though $n_{i}$ play no role in any of the identified circuits.
@@ 3.1 Identifying Circuits
We next focused on different circuits involving these edges. We considered both general edge sets of types (1) and (2), and the local subset ($x_{i}\rightarrow x_{i+1}$ and $y_{i}\rightarrow y_{i+1}$); the latter were sufficient in all but one case ($y_{i}\rightarrow y_{j}$ needed for Capitalization). For each task, we chose the simplest circuit achieving $\geq 0.9$ at $N=3$, or (if there is none) the circuit achieving highest accuracy at $N=3,10$, all based on the 2B model’s accuracies (Tables 2). We illustrate the result for Country-Capital in Appendix, Figure 1; circuits for all five tasks are shown in Appendix, Figure 24. At 2B parameters and 10 few-shots, the circuits explained at least 90% of the full model’s performance, far exceeding the aggregation-only circuit on some tasks (Figure 2). We next evaluated the circuits on Gemma-2 9B and Gemma-2 27B, using the same methodology (Tables 5, 6). At 10 shots, the circuits that we had chosen based on the 2B model explained at least 90% of the full model’s performance in these model. Importantly, the Aggregation circuit again underperformed the circuits involving contextualization on most tasks (Figure 2).
@@
In order to identify the heads responsible for the information flow we have identified, we next projected the position-level circuits to activation-level circuits. Here, nodes are activations of attention heads indexed by both their positions and layers; edges are directed arcs between the input and output of individual attention heads. More precisely, heads can be defined as tuples of (layer, head, position), and edges can be described as (layer, head, start position, end position). We used a gradient-based method (Michel et al., 2019; Syed et al., 2023) (see Appendix F for technical details) to obtain a circuit on the level of activations with edges defined by attention heads, in the 3-shot setup (Figures 913). The circuit reveals that the contextualization step tends to precede the aggregation step: contextualization happens in the lower half layers, most aggregation edges are in the middle and upper layers. Recall that prior work has established the existence of function vector heads, specific heads aggregating causally relevant task information at $t_{N+1}$ (Section 2). We next identified function vector heads by using the method described in Todd et al. (2024) for each task independently and selecting the top-10-scoring heads for each task. These heads were largely part of the identified circuits and mostly contributed to the $y_{i}\rightarrow t_{N+1}$ edges (Appendix, Figure 26). This confirms that the circuits we have identified describe the buildup of task information up to the function vector heads.
Figure 3: Patching experiments manipulating the information about the input available to each edge in the full circuit, at the example of the Country-Capital task, 10-shot. For each ablation, we indicate whether the drop in performance is statistically significant compared to the full circuit performance according to a binomial test with $\alpha=0.05$. See Figure 8 for results from all tasks.
@@ 3.2 Which Information is Routed?
We vary the information about the input that is available to a particular edge by constructing a contrastive input sample on which certain aspects of the information have been altered (Appendix H, Figure 22), and patching the K and V activations directly feeding into the edge with the corresponding activations computed on that contrastive input (Appendix, Figure 28). For instance, when patching an edge $y_{i}\rightarrow t_{N+1}$, we modify the key and value activations associated with the positions of $y_{i}$ across all attention heads in the model. These modifications occur specifically when the activations are queried from the positions corresponding to $t_{N+1}$. This patching is done in parallel in each layer and head. All patching is done within the overall circuit identified for the task, i.e., edges not present in the circuit are ablated, and paths not present in the overall circuit play no role even in the patched version. For instance, to determine whether the edges $y_{i}\rightarrow t_{N+1}$ transport contextualized information about preceding few-shots $(x_{i^{\prime}},y_{i^{\prime}})$ ($i^{\prime}<i$), we patch all inputs to this edge at position $y_{i}$ with activations computed on an input where the preceding few-shots have been corrupted; by varying the degree of corruption, we analyze which specific kinds of information are transported (Appendix H, Figure 22). We note that this methodology allows us to focus on causally relevant information used by downstream components. It is thus fundamentally different from supervised probing, which does not distinguish between information used or unused by downstream components. We focused on the 2B model for tractability.
@@
We next moved to analyzing the information transported by the contextualizing edges, starting with the edges connecting $y$’s ($y_{i}\rightarrow y_{i+1}$ or general $y_{i}\rightarrow y_{j}$). We found that these edges transport both information about $f$ and about $\mathcal{Y}$. Patching the $y_{i}\rightarrow y_{i+1}$ edges for Country-Capital only decreases accuracy significantly when $y$ leaves $\mathcal{Y}$, but not under other corruptions (Figure 3). In Capitalization, at 3 shots, replacing with $y\in\mathcal{Y}$ hurts much less than a general ablation of $y$ (accuracy 81% vs 56 %; Appendix F, Figure 8), which is as harmful as full corruption of these edges and leads to a large increase in copying behavior; nonetheless, accuracy still drops compared to ablations keeping $(x,y)\in f$. At 10 shots, Capitalization only responds to ablations where $y$ leaves $\mathcal{Y}$. In both cases, ablating the output space information leads to an increased number of copy-type mistakes (Appendix H, Figure 19).
As noted before, edges do not provide specific token information used by downstream components at $t_{N}$: we verified that a token mismatch created by intervening on $(x_{i-1},y_{i-1})$ for all $y_{i}\rightarrow y_{j}$ edges but not the $y_{i}\rightarrow t_{N+1}$ edges has no discernible effect on accuracy if the functional relation between is preserved by the prompt used for intervention.
Overall, our causal interventions warrant the conclusion:
@@ 3.3 A Contextualization subcircuit is important in the presence of ambiguity
We conducted a controlled experiment aiming to stress-test the aggregation step. We designed the following hard version of the Present-Past task. We take advantage of the fact that some English verbs have identical present and past tense forms (e.g., “put”, “spread”). In a prompt for the Present-Past task, examples using such verbs are thus ambiguous between the Present-Past task and a simpler Copying task (Appendix H, Figure 23.1). In this experiment, most individual examples do not give full task information, and may in fact provide misleading information. We hypothesize that direct aggregation of uncontextualized per few-shot representations as hypothesized by prior work may be less useful in this case, because individual examples might, on their own, provide conflicting task information. The contextualization step might help resolve this ambiguity before the aggregation step.
Even when 7 out of 10 few-shot examples are ambiguous, the full model continues to perform the task at very high accuracy (95%). Whereas the aggregation (Aggregation)-only circuit had performed at 93% on the standard Present-Past task (and similarly on the Copying task), its performance on the ambiguous prompts dropped to 56% (Appendix H, Table 9).
Figure 5: Patching with ambiguous prompts (Section 3.3): Patching $y_{i}\rightarrow t_{N+1}$ connections on only ambiguous examples (bottom) using activations taken on all-ambiguous (red) or all-unambiguous (green) prompts leads to no significant drop. In contrast, performing these patches on only unambiguous examples (top) leads to a drastic drop, even when patching with an all-unambiguous prompt (green).
@@
We investigated two hypotheses as to how contextualization might resolve ambiguity (Figure 4):
@@
In contrast, patching the unambiguous examples had great effect: Accuracy dropped to almost zero when patching with ambiguous prompts (in which case, all task information would be routed through the ambiguous examples, but we saw that they do not causally provide disambiguation signal via $y_{i}\rightarrow t_{N+1}$). Strikingly, accuracy dropped drastically even when patching with fully unambiguous prompts, showing that contextualizing the unambiguous examples with awareness of the presence of ambiguous examples is key to downstream performance. Overall, our results allow us to reject H1. They are as expected under H2:

### v4

@@ 2 Background: Aggregation and Function Vectors
Prior work on the mechanics of ICL in LLMs provides a few robust insights. Multiple studies document the existence of task vectors or function vectors (Hendel et al., 2023; Todd et al., 2024; Song et al., 2025; Kharlapenko et al., 2025; Yin and Steinhardt, 2025): Several attention heads at the last prompt token (in our case, $t_{N+1}$) output vectors that, across ICL tasks, encode task information. Erasing them prevents ICL; patching them elsewhere leads to execution of the task. Todd et al. (2024) specifically operationalize Function Vector Heads as heads whose activations, when patched from valid ICL prompts to shuffled uninformative prompts, can most increase the likelihood of the desired target. Yin and Steinhardt (2025) find that Function Vector Heads are more important to ICL than Induction Heads, hypothesized to be key in prior work (Olsson et al., 2022; Crosbie and Shutova, 2024) (cf. also Bansal et al. (2023)). However, it remains largely open what circuit assembles this information, and how information is assembled from the different few-shot examples to these function vectors.
@@
Kharlapenko et al. (2025) discover a sparse feature circuit in Gemma-2 2B, using sparse autoencoder (SAE) features encoding task information (in few-shot examples) and features that execute them (at $t_{N+1}$). However, the circuit does not distinguish between inputs and outputs for different few-shot examples, leaving open whether task information is created independently at each example or whether the representations are dependent across few-shot examples. Cho et al. (2025) suggest a different circuit: (1) assembling an encoding of each few-shot example’s input and moving it to few-shot label position, (2) attending to labels whose input is similar to the query, and copying these labels to the query. Their experiments are restricted to semantic text classification (e.g., sentiment), in which few-shot inputs typically are either semantically aligned or opposed. In addition, the arguments are largely based on evaluating representational similarity and on supervised probing, which does not establish that this mechanism is causally sufficient. While a causal intervention, through zeroing out the relevant components, shows this mechanism is relevant to ICL (Section 5.1 in Cho et al. (2025)), it leaves open if the mechanism alone can fully explain the behavior, or whether it is only a part of the overall mechanism. Overall, while function vectors play an important role for many ICL tasks, the information flow leading to aggregation of task information across few-shot examples remains unclear. To foreshadow our results, we show that, while parallel aggregation of per few-shot representations is important, the overall circuit is task-dependent, and may require a variety of additional components which, as we will show, contextualize the per few-shot representations on other few-shot examples.
@@ 3 Results
We focus on Gemma-2, especially its 2B version, but our methodology allows us to also investigate larger variants (9B and 27B). We evaluated the model on tasks from Todd et al. (2024), and determined five tasks on which the 2B model achieves high (between $88\%$ and $100\%$) accuracy: Copying  [^2] This is not included in Todd et al. (2024), but we added it due to its naturalness., Present-Past, Capitalization, Country-Capital, Person-Sport (Figure 2). [^3] We also evaluated six other models at the 2B/3B scale, finding that they overall underperformed Gemma-2 2B at 10 shots on these tasks (Table 7), further motivating our focus on Gemma-2. We decided to focus on tasks with high accuracy in order to obtain a clear signal, as for these tasks there is a clear understanding of the input-output behavior exhibited by the full model. We focused experiments on shorter (3-shot) and longer (10-shot) prompts.
[…]
3.1 Identifying Circuits
To localize the behavior of the model, we use patching, an approach widely adopted in the literature (Wang et al., 2022; Hanna et al., 2023). With this technique, a model is viewed as a computation graph with activations in different layers as nodes, and computations between them as edges. We then ablate some of the edges in the graph, forcibly replacing computation along a specific edge with the one computed on the counterfactual input. Counterfactual inputs are designed to erase relevant information from the input while leaving other information intact for localization of model behavior. For example, when substituting "Berlin" with counterfactual "Paris", we isolate computations specific to the city’s identity, not those activated by its category (city) or token-type (word). If ablating a set of edges does not lead to a drop in the model’s performance, then the information unique to the original input relative to the counterfactual input, which was transferred along these edges, is not causal for the model’s prediction. This allows to discover circuits – subgraphs of the full model’s computation graph explaining a major part of model performance on a specific task. In contrast to the standard approach, we focus on the information flow between positions in a prompt; consequently, we differentiate between activations of the same heads at different positions. We keep nodes for $x_{i}$ and $y_{i}$ in different examples distinct, which allows us to analyze the information flow between few-shot examples. [^4] This contrasts with Kharlapenko et al. (2025), who collapse all $x_{i}$ and $y_{i}$ nodes into one node each.
We first discover a position-level circuit. Nodes are the positions in the prompt; edges are directed arcs between them, representing the flow of information (Figure 1). Importantly, in position-level circuits we collapse all edges across layers into a single edge. Ablating such an edge means ablating all corresponding edges in every attention head and layer. For illustration, $x_{2}$ in a position-level circuit can only receive four incoming edges: from $x_{1},t_{1},y_{1}$, or $n_{1}$.  [^5] The edges from/to bos/eos tokens are always ablated (Appendix E). Position-level circuits provide useful and interpretable upper bounds on information flow, though they do not distinguish between computations happening in different layers. We treated all tokens within a few-shot input $x_{i}$ or output $y_{i}$ together, as the number of tokens varies between few-shot examples. As positions, we distinguish the few-shot inputs $x_{i}$ and outputs $y_{i}$ and the separators within ($t_{i}$) and between ($n_{i}$) examples; though $n_{i}$ play no role in any of the identified circuits.
@@ 3.1 Identifying Circuits
We next focused on different circuits involving these edges. We considered both general edge sets of types (1) and (2), and the local subset ($x_{i}\rightarrow x_{i+1}$ and $y_{i}\rightarrow y_{i+1}$); the latter were sufficient in all but one case ($y_{i}\rightarrow y_{j}$ needed for Capitalization). For each task, we chose the simplest circuit achieving $\geq 0.9$ at $N=3$, or (if there is none) the circuit achieving highest accuracy at $N=3,10$, all based on the 2B model’s accuracies (Tables 2). We illustrate the result for Country-Capital in Appendix, Figure 1; circuits for all five tasks are shown in Appendix, Figure 24. At 2B parameters and 10 few-shots, the circuits explained $\geq 90\%$ of the full model’s performance, far exceeding the aggregation-only circuit on some tasks (Figure 2). We next evaluated the circuits at 9B and 27B, using the same methodology (Tables 5, 6). At 10 shots, the circuits that we had chosen based on the 2B model explained at least 90% of the full model’s performance in these model. Importantly, the Aggregation circuit again underperformed the circuits involving contextualization on most tasks (Figure 2).
@@
In order to identify the heads responsible for the information flow we have identified, we next projected the position-level circuits to activation-level circuits. Here, nodes are activations of attention heads indexed by both their positions and layers; edges are directed arcs between the input and output of individual attention heads. More precisely, heads can be defined as tuples of (layer, head, position), and edges can be described as (layer, head, start position, end position). We used a gradient-based method (Michel et al., 2019; Syed et al., 2023) (see Appendix F for technical details) to obtain a circuit on the level of activations with edges defined by attention heads, in the 3-shot setup (Figures 913). The circuit reveals that the contextualization step tends to precede the aggregation step: contextualization happens in the lower half layers, most aggregation edges are in the middle and upper layers. Prior work has established the existence of function vector heads, specific heads aggregating causally relevant task information at $t_{N+1}$ (Section 2). We next identified function vector heads by using the method described in Todd et al. (2024) for each task independently and selecting the top-10-scoring heads for each task. These heads were largely part of the identified circuits and mostly contributed to the $y_{i}\rightarrow t_{N+1}$ edges (Appendix, Figure 26). This confirms that the circuits we have identified describe the buildup of task information up to the function vector heads.
Figure 3: Patching experiments manipulating information about the input available to each edge in the full circuit, at the example of the Country-Capital task, 10-shot. For each ablation, we indicate whether the drop in performance is statistically significant compared to the full circuit performance in a binomial test with $\alpha=0.05$. See Figure 8 for results from all tasks.
@@ 3.2 Which Information is Routed?
We vary the information about the input available to a particular edge by constructing a contrastive input sample in which certain aspects have been altered (Appendix H, Figure 22), and patching the K and V activations directly feeding into the edge with the corresponding activations computed on that contrastive input (Appendix, Figure 28). For instance, when patching an edge $y_{i}\rightarrow t_{N+1}$, we modify the key and value activations associated with the positions of $y_{i}$ across all attention heads in the model. These modifications occur specifically when the activations are queried from the positions corresponding to $t_{N+1}$. Patching is done in parallel in each layer and head. All patching is done within the overall circuit identified for the task, i.e., edges not present in the circuit are ablated, and paths not present in the overall circuit play no role even in the patched version. For instance, to determine whether the edges $y_{i}\rightarrow t_{N+1}$ transport contextualized information about preceding few-shots $(x_{i^{\prime}},y_{i^{\prime}})$ ($i^{\prime}<i$), we patch all inputs to this edge at position $y_{i}$ with activations computed on an input where the preceding few-shots have been corrupted; by varying the degree of corruption, we analyze which specific information is transported (Appendix H, Figure 22). We note that this methodology allows us to focus on causally relevant information used by downstream components. It is thus fundamentally different from supervised probing, which does not distinguish between information used or unused by downstream components. We focus on the 2B model for tractability.
@@
We next moved to analyzing the information transported by the contextualizing edges, starting with the edges connecting $y$’s ($y_{i}\rightarrow y_{i+1}$ or general $y_{i}\rightarrow y_{j}$). We found that these edges transport both information about $f$ and about $\mathcal{Y}$. Patching the $y_{i}\rightarrow y_{i+1}$ edges for Country-Capital only decreases accuracy significantly when $y$ leaves $\mathcal{Y}$, but not under other corruptions (Figure 3). In Capitalization, at 3 shots, replacing with $y\in\mathcal{Y}$ hurts much less than a general ablation of $y$ (accuracy 81% vs 56 %; Appendix F, Figure 8), which is as harmful as full corruption of these edges and leads to a large increase in copying behavior; nonetheless, accuracy still drops compared to ablations keeping $(x,y)\in f$. At 10 shots, Capitalization only responds to ablations where $y$ leaves $\mathcal{Y}$. In both cases, ablating the output space information leads to an increased number of copy-type mistakes (Appendix H, Figure 19). As noted before, edges do not provide specific token information used by downstream components at $t_{N}$: we verified that a token mismatch created by intervening on $(x_{i-1},y_{i-1})$ for all $y_{i}\rightarrow y_{j}$ edges but not the $y_{i}\rightarrow t_{N+1}$ edges has no discernible effect on accuracy if the functional relation between is preserved by the prompt used for intervention. Overall, our causal interventions warrant the conclusion:
@@ 3.3 A Contextualization subcircuit is important in the presence of ambiguity
We conducted a controlled experiment aiming to stress-test the aggregation step. We designed the following hard version of the Present-Past task. We take advantage of the fact that some English verbs have identical present and past tense forms (e.g., “put”, “spread”). In a prompt for the Present-Past task, examples using such verbs are thus ambiguous between the Present-Past task and a simpler Copying task (Appendix H, Figure 23.1). In this experiment, most individual examples do not give full task information, and may in fact provide misleading information. We hypothesize that direct aggregation of uncontextualized per few-shot representations as hypothesized by prior work may be less useful in this case, because individual examples might, on their own, provide conflicting task information. The contextualization step might help resolve this ambiguity before the aggregation step. Even when 7 out of 10 few-shot examples are ambiguous, the full model continues to perform the task at very high accuracy (95%). Whereas the aggregation (Aggregation)-only circuit had performed at 93% on the standard Present-Past task (and similarly on the Copying task), its performance on the ambiguous prompts dropped to 56% (Appendix H, Table 9).
@@
We investigated two hypotheses as to how contextualization resolves ambiguity (Figure 4):
@@
Figure 5: Patching $y_{i}\rightarrow t_{N+1}$ connections on only ambiguous examples (bottom) using activations taken on all-ambiguous (red) or all-unambiguous (green) barely hurts. In contrast, performing these patches on only unambiguous examples (top) leads to a drastic drop, even when patching with an all-unambiguous prompt (green).
@@
In contrast, patching the unambiguous examples had great effect: Accuracy dropped to almost zero when patching with ambiguous prompts (in which case all information would be routed through the ambiguous examples, but we saw that they do not causally provide disambiguation signal via $y_{i}\rightarrow t_{N+1}$). Strikingly, accuracy dropped drastically even when patching with fully unambiguous prompts, showing that contextualizing the unambiguous examples with awareness of the presence of ambiguous examples is key to downstream performance. Overall, our results allow us to reject H1. They are as expected under H2: